OpenStudy (anonymous):

Find an equation of the line that is tangent to the graph of f and parallel to the given line. Function f(x) = x^2 − x Line x + 3y − 6 = 0

OpenStudy (goformit100):



OpenStudy (stamp):

\[\]\[f(x)=x^2-x\]\[y=(6-x)/3\]\[m=-1/3\]\[f'(x)=2x-1\]\[-1/3=2x-1\]\[2/3=2x\]\[x=1/3\]\[f'(1/3)=-1/3\]\[y=-x/3+b\]\[(1/3,\ f(1/3))=(1/3,\ -2/9)\]\[b=y+x/3\]\[b=-2/9+1/9\]\[b=-1/9\]\[line=-x/3-1/9\]